• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ole Hansen

    Log entry time 13:39:47 on July 08, 2010

    Entry number 330807

    This entry is a followup to: 330677


    keyword=L-arm VDC wiremaps - u1 fixed, minor u2 problem

    Here are the VDC wiremaps for run 1852, cut on event type 5 (clean L-arm single arm electron trigger - I think) and time. The problem with two bad cards in u1 is apparently gone, however, as already observed by the owl shift, there is one dead card in u2 now (wires 336-351, 2nd card from low momentum end).

    The spectra have too much structure overall, indicating noise and/or crosstalk issues with the high-gain amplifier cards in general.

    Also, there is a very clear dip in all spectra towards low momentum (high wire numbers). This dip was not present during yesterday's day shift. The structure in this region changes significantly for other trigger types - it is a small dip followed by a large peak for trigger type 4. Something odd seems to be going on, especially since no changes seem to have been made since yesterday's day and this morning's owl shift.

    A copy of this log entry has been emailed to: bogdanw, kalyan

    Figure 1

    Figure 2

    Figure 3

    Figure 4